# Enhancing Socially-Aware Robot Navigation through Bidirectional Natural Language Conversation

### v1

@@ IV-C1 Discrete Action Space Dimension
IV-C2 Effectiveness of LLM module
In this section, we validate the effectiveness of the LLM component within the HSAC-LLM model. To this end, we conducted a comparative experiment by removing the LLM module responsible for natural language bidirectional interaction with pedestrians from the HSAC-LLM model. Specifically, when the robot detected a potential collision risk, it merely issued an auditory warning to pedestrians without the capability to receive vocal feedback from them. Table III presents a comparison between the HSAC model, with the LLM module removed, and the original HSAC-LLM model across three scenarios. The results from the table show that with the incorporation of the LLM module, there was an increase in success rates and a decrease in both collision rates and the number of steps to reach the target across all three scenarios. Notably, the improvement in collision rates was more significant, especially in the Square and Crosswalk scenarios, with reductions of 42.86% and 30.77%, respectively. This demonstrates that the integration of LLM enables the robot to better understand natural language and communicate with pedestrians, thereby eliminating collisions that could arise from inadequate communication. The inclusion of the LLM module not only markedly improves navigational safety by minimizing potential collisions but also significantly enhances the robot’s capability in social navigation. By enabling more nuanced natural language communication with pedestrians, this development highlights the essential role of advanced language processing in enriching human-robot interactions within dynamic environments. This progress in social navigation illustrates the substantial potential of integrating sophisticated language models into robotic systems, paving the way for more intuitive and effective human-robot collaborations, especially in complex social settings.
@@ V CONCLUSIONS
This paper introduces HSAC-LLM, a novel approach to socially-aware navigation that combines deep reinforcement learning with large language models. This approach enables bidirectional interactions based on natural language between robots and pedestrians. Experiments in 2D simulations, the Gazebo environment, and real-world settings further demonstrate the superior performance of our model in navigation and obstacle avoidance. We believe this innovative paradigm not only enhances the safety and efficiency of robot navigation but also facilitates harmonious coexistence between humans and robots in complex social environments.

### v2

@@ IV-C Experiments in Real-World environment
IV-D Ablation Study
IV-D1 Discrete Action Space Dimension
To assess how discrete action space dimensionality affects navigation performance, we tested HSAC-LLM with both 2-D ($\bm{a}_{c}\in\{\text{none},\text{stop}\}$) and 4-D discrete action spaces across Square, Hallway, and Crosswalk environments. Table II presents the results based on Success Rate, Collision Rate, Steps Taken, and Training Time metrics. The 4-D discrete action space consistently outperformed the 2-D configuration across all environments. In the Square environment, success rate improved from 0.90 $\pm$ 0.029 to 0.91 $\pm$ 0.022, while collision rate decreased from 0.10 $\pm$ 0.029 to 0.08 $\pm$ 0.022. Similar improvements occurred in the Hallway scenario (success rate: 0.92 $\pm$ 0.036 to 0.93 $\pm$ 0.013; collision rate: 0.08 $\pm$ 0.036 to 0.07 $\pm$ 0.013). The Crosswalk environment showed the most significant improvement, with success rate increasing from 0.87 $\pm$ 0.026 to 0.91 $\pm$ 0.022 and collision rate decreasing from 0.13 $\pm$ 0.026 to 0.09 $\pm$ 0.022. While the 4-D configuration required slightly longer training times (e.g., increasing from 1.67 to 1.89 in Hallway and from 1.78 to 2.04 in Crosswalk), the performance gains justify this trade-off. These results demonstrate that expanding the discrete action space enhances HSAC-LLM’s ability to navigate complex environments safely and efficiently, particularly in scenarios requiring nuanced interactions with pedestrians.
@@ IV-D1 Discrete Action Space Dimension
IV-D2 Effectiveness of LLM module
To evaluate the LLM component’s effectiveness in HSAC-LLM, we conducted experiments comparing the full model against a version with the LLM module removed (HSAC). Without the LLM module, the robot could only issue warnings to pedestrians but couldn’t process their verbal feedback. Table III compares both models across three scenarios. Results show that incorporating the LLM module improved success rates and reduced collision rates and navigation steps in all scenarios. Collision rate improvements were particularly significant in Square and Crosswalk environments, with reductions of 42.86% and 30.77%, respectively. The LLM integration enables better natural language understanding and communication with pedestrians, preventing collisions caused by communication gaps. This enhancement not only improves navigational safety but also significantly advances the robot’s social navigation capabilities, demonstrating the value of sophisticated language processing in human-robot interactions within dynamic environments.
@@ V CONCLUSIONS
This paper presented HSAC-LLM, a novel framework integrating deep reinforcement learning with large language models for socially-aware robot navigation. Our approach enables natural language interactions between robots and humans, enhancing the robot’s ability to understand human intentions in dynamic environments. HSAC-LLM models both continuous navigation actions and discrete conversation actions. Experiments in 2D simulations, Gazebo, and real-world settings demonstrate superior performance in safety, efficiency, and human comfort compared to state-of-the-art DRL methods. This work advances intelligent human-robot interactions in social environments. Future work will focus on improving language understanding, real-time adaptation to complex scenarios, and large-scale real-world deployment.